2. Let и(x, y, 2) =y+y23, t = rse, y = rse- = r’s sint Compute ди ди at the point r = 2, 8= 1, t= 0 дѕ де ət ди
Answer #1

Given verset, zarstet and -z =rs sint. t se Therz, da ar Әд ar centaret, perset sket 34 as erset 22 h - -rsket = 2rse - az aSo, As 112,2,2) = 248 +2 =23 au 4X4 08 au 3 24+ 22² 031 du az = 321 2,21=2, so at Y=2, S=1,t=o that t=o that is at 3= we 2 =By chain rule, Әи Əs au ат дчӘд Оч dz त्र Əs тәу әѕ təri əs (64 x 2) + (16х4) (14х4) + (0 хо) 12.8 + 4 + 0 — 192. Әu år au да